hello I need help solving this question,
if a goat is tethered to the outside corner of a barn by 16 foot rope, what that is the maximum area of grass the goat can eat? (the barn forms a right angle).

thank you!!

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

Depends on the size of the barn.

Answer by stanbon(75887)

if a goat is tethered to the outside corner of a barn by 16 foot rope, what that is the maximum area of grass the goat can eat? (the barn forms a right angle).
----
Assuming the sides of the barn are <= 16 ft, the
maximum feeding area is (3/4)th of a circle with diameter = 16 ft
----
Area = (3/4)pi*8^2 = 48pi sqft
----------------
Cheers,
Stan H.
-----------